1. In the speed of sound experiment, the length of the tube was 93 cm and the time between the generated and detected sound was measured to be 6.4 ms. Determine the distance (in m) traveled by sound in air from the given data.

2.In the speed of sound experiment, the length of the tube was 101 cm and the time between the generated and detected sound was measured to be 5.8 ms. Determine the speed of sound in air from the given data.
Answer #1

Solution :

Given :

Length of tube (L) = 93 cm = 0.93 m

and the time between the generated and detected sound was measured to be 6.4 ms

Thus : t = 6.4 ms = 0.0064 s

For the sound to be detected, it has to travel twice the length of the tube. Once while going in and once more after reflecting back.

So, the distance traveled is = 2 L = 2(0.93 m) = 1.86 m.

And, the speed of sound is: V = 2L / t = (1.86 m) / (0.0064 s) = 290.625 m/s.